OpenStudy (anonymous):

I need help simplifying this problem. 4y^3/3bk^2 - 5bx/6d^3k I got 1xy^3/3bd^3k but I think that is wrong.

OpenStudy (campbell_st):

well you need a common denominator of 6bd^3k^2 so the 1st term needs the numerator multiplied by 2d^3 and the 2nd term needs the numerator multiplied by bk so you'll have \[\frac{4y^3 \times 2d^3}{6bd^3k^2} - \frac{5bx \times bk}{6bd^3k^2}\]

OpenStudy (anonymous):

\[\large{{4y^3 \over 3bk^2} \color{red}{\frac{2d^3}{2d^3}}- {5bx \over 6d^3k}\color{red}{\frac{bk}{bk}}}\]\[\large \frac{8y^3d^3-5xb^2k}{6bk^2d^3}\]

OpenStudy (anonymous):

Thank you both of you all!

OpenStudy (anonymous):

@ChmE do the variables have to be in order by alphabet or it doesn't matter?

OpenStudy (campbell_st):

it doesn't matter, the normal convention for algebra is in order but you can't be marked wrong if its not

OpenStudy (anonymous):

Thank you so much @campbell_st